### v1

# Constrained Language Generation with Discrete Diffusion Models

Abstract
Constraints are critical in text generation as LLM outputs are often unreliable when it comes to ensuring generated outputs adhere to user defined instruction or general safety guidelines. To address this gap, we present Constrained Discrete Diffusion (CDD), a novel method for enforcing constraints on natural language by integrating discrete diffusion models with differentiable optimization. Unlike conventional text generators, which often rely on post-hoc filtering or model retraining for controllable generation, we propose imposing constraints directly into the discrete diffusion sampling process. We illustrate how this technique can be applied to satisfy a variety of natural language constraints, including (i) toxicity mitigation by preventing harmful content from emerging, (ii) character and sequence level lexical constraints, and (iii) novel molecule sequence generation with specific property adherence. Experimental results show that our constraint-aware procedure achieves high fidelity in meeting these requirements while preserving fluency and semantic coherence, outperforming auto-regressive and existing discrete diffusion approaches.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable advancements in reasoning, enabling progress across diverse domains. Beyond traditional NLP applications, LLMs are increasingly being integrated into decision-making pipelines in embodied agents, autonomous scientific systems, and industrial automation [1, 2, 3, 4]. Despite this promise, the stochastic nature of LLM outputs makes these models untrustworthy when required to adhere to even simple logical constraints, domain-specific knowledge, and safety-critical requirements [5].
This limitation manifests in well-documented issues such as hallucinations, misinformation, and toxic outputs, presenting severe risks in real-world deployments [6, 7]. Furthermore, generative language models remain susceptible to producing harmful or biased text, or even disregarding user instructions, particularly when confronted with adversarial or ambiguous prompts [8]. The inability of generative models to adhere to strict operational and safety constraints severely undermines their reliability in scientific and industrial applications. For instance, in autonomous laboratories, LLM-driven systems may misrepresent experimental constraints or respond to adversarial triggers, leading to erroneous instructions for synthesizing hazardous or unstable compounds [9, 10]. This lack of assurance raises concerns and diminishes trust in LLM-driven systems, limiting their adoption in safety-critical environments.
Recent efforts have attempted to improve constraint satisfaction in LLM outputs through soft alignment methods via reinforcement learning from human feedback (RLHF) [11], knowledge injection [12, 13], or heuristic filtering [14]. However, while these approaches reduce harmful outputs, they lack mechanisms to consistently provide conformance with domain-specific or complex logical requirements. Furthermore, techniques including post-processing and rejection sampling, either completely discard the generated content or significantly alter it, thus compromising the creative quality of the original generation.
Figure 1: Comparison of Constrained Discrete Diffusion (CDD) and baseline models. CDD imposes constraints without sacrificing fluency or expressiveness.
To address these challenges, we propose integrating token generation with a novel differentiable optimization process that directly guides outputs to meet the imposed constraints. A key enabling component is the recent development of discrete diffusion models [15, 16, 17, 18], which offer an alternative to conventional auto-regressive architectures and have demonstrated strong performance across language generation tasks [4, 19]. Analogous to continuous diffusion models, discrete diffusion progressively refines corrupted sequences into coherent outputs. Notably, these models have achieved perplexity scores comparable to similarly-sized auto-regressive architectures on recent benchmarks [19, 16]. In this work we show how combining discrete diffusion with structured constraints during the sampling process enables the imposition of domain or safety requirements at every denoising step, without sacrificing model creativity or expressiveness.
Our approach is tested with three representative use cases: (1) toxicity mitigation with tunable thresholds for safe LLM deployment, (2) instruction-following tasks requiring precise symbolic constraints, and (3) property adherence for molecular sequence generation. Experimental results demonstrate our method’s ability to obtain 0% threshold violations for toxicity, 100% satisfaction for instruction following constraints, and up to 3x valid and novel molecule generations.
Contributions. Our contributions are as follows:
  1. 1.
    We introduce CDD, a novel framework for constrained language generation that integrates discrete diffusion models with differentiable optimization, enforcing constraints directly within the diffusion sampling process.
  2. 2.
    To enable the constrained optimization process during the token generation, we formulate a projection based operator, which we implement using a Lagrangian Dual based framework.
  3. 3.
    We analyze CDD performance on a set of complex natural language and structured text generation tasks, reporting state-of-the-art results for both constraint satisfaction and qualitative metrics.
2 Related Work
Recent efforts in controllable text generation have largely focused on constrained decoding, where the output is dynamically restricted to adhere to predetermined syntactic, semantic, or lexical rules. Approaches in this area modify the decoding process to prune the vocabulary, ensuring that only tokens compliant with a formal grammar are considered [20], or employ external models to filter outputs when direct access to the model’s logits is limited [21]. Other methods have also adapted beam search to encourage the presence or absence of specific words [22, 23]. Although these techniques effectively guide the generation process, they depend on augmented sampling heuristics which encourage, but frequently fail to provide, satisfaction of even simple constraint sets. Therefore, a recent line of research has attempted at mapping natural language into formal logic representations to verify factual and logical consistency [24, 25, 26], and by enforcing feasibility constraints with external modules [27, 28]. However, these approaches, while effective in their respective domains, often require additional, domain-specific components to ensure that the generated outputs are not only syntactically valid but also adhere to practical or logical requirements. Most similar to our work, Kumar et al. [29] and Dathathri et al. [30] introduce gradient-based corrections into non-autoregressive and autoregressive language models, respectively, to steer text generation towards particular attributes; however, these methods suffer from notable short-comings, as outputs from Kumar et al. [29] exhibit poor performance (e.g., higher perplexity), and outputs from Dathathri et al. [30] often still violate imposed specifications. Our work draws on these insights, seeking to embed both syntactic and semantic constraints directly within the generative process for enhanced control and reliability.
3 Preliminaries
Generative Diffusion Models.
Generative diffusion models [31, 32] have revolutionized high-fidelity creation of complex data and serve as key components in state-of-the-art image and video synthesis pipelines [33, 34, 35]. Formally, they operate in two phases: a forward phase, which gradually introduces noise to clean data samples, followed by a reverse denoising phase, in which a deep neural network is trained to iteratively remove noise. The forward process defines a Markov chain $\{\bm{x}_{t}\}_{t=0}^{T}$, with initial sample $\bm{x}_{0}\sim p(\bm{x}_{0})$ and each transition $q(\bm{x}_{t}|\bm{x}_{t-1})$ obtained by adding Gaussian noise that increases as $t\to 1$. The reverse process relies on a learned network that directly predicts the noise added at each step, enabling the gradual recovery of the original data distribution. Once trained, this network can then be used to iteratively “denoise” random noise samples back into data-like samples. This step is also known as sampling.
Discrete Diffusion Language Models.
While diffusion models were originally developed for continuous data, they have recently been extended to discrete domains such as text, enabling non-autoregressive generation [36, 18, 16, 37]. In contrast to auto-regressive models which predict tokens one by one, discrete diffusion methods generate entire sequences in parallel by first corrupting sequences through a forward noising process and then iteratively reconstructing them with a learned reverse process. This is a key enabler recognized by this work, which exploits this modus operandi to impose global constraints while simultaneously maintaining high fidelity.
Let $\bm{x}_{0}=(\bm{x}_{0}^{1},\ldots,\bm{x}_{0}^{L})$ denote an input sequence of size $L$, where each token $\bm{x}_{0}^{i}\in\mathbb{R}^{V}$ is represented as a one-hot vector over a vocabulary of size $V$. In discrete diffusion models, the forward process produces, at time $t\in[0,T]$, a sequence of values $\bm{x}_{t}\in\mathbb{R}^{L\times V}$ that parametrizes probability distributions for each token $\bm{x}_{t}^{i}$ in the sequence. We denote the corresponding sequence of predicted tokens by $\bm{x}^{\star}_{t}=\arg\max(\bm{x}_{t})$ where the $\arg\max$ operator is applied to every member $\bm{x}_{t}^{i}$ of the sequence $\bm{x}_{t}$. The diffusion process specifies a forward transition, defined as:
${q}(\bm{x}_{t}\mid\bm{x}_{0})=\mathrm{Cat}\bigl{(}\bm{x}_{t};\alpha_{t}\bm{x}_{0}+(1-\alpha_{t})\,\nu\bigr{)},$ (1)
where $\alpha_{t}$ is a schedule that decreases over time, $\nu$ is a fixed reference distribution that specifies the type of corruption applied to each token, and $\mathrm{Cat}(\cdot;p)$ is the categorical distribution over probability vector $p\in\Delta^{V}$. For example, the Masked Diffusion Language Model (MDLM) [16] uses a one-hot vector corresponding to the [MASK] token. Here, once a token is replaced by [MASK] (i.e., absorbed), it remains fixed during the forward process, and once unmasked during the denoising process it remains unchanged. In contrast, Uniform Diffusion Language Models (UDLM) [17], uses $\nu$ as the uniform distribution over the vocabulary. Conversely, this instantiation allows tokens to be re-perturbed in later time steps.
Figure 2: Illustration of CDD’s projection step embedded throughout the sampling process.
4 Constraint-aware Language Generation
This section introduces our key contribution: a paradigm to provide constraint adherence during the token generation process. Our novel method, Constrained Discrete Diffusion (CDD) alternates differentiable projections with iterative denoising steps during the sampling phase, ensuring that generated sequences satisfy predefined rules.
4.1 Projected Diffusion Processes
Discrete diffusion models generate samples through a stochastic denoising process, wherein a sequence of discrete states $\bm{x}_{t}$ evolves toward a target distribution $p(\bm{x}_{0})$. At each iteration, the model updates the token representation via a transition function $\mathcal{U}_{\theta}$, parametrized by a learned weight vector $\theta$, producing the next-step estimate:
$\bm{x}_{s}=\mathcal{U}_{\theta}(\bm{x}_{t})$
where $s$ is the subsequent timestep in the continuous time reverse process and $0\leq s<t\leq T$. Here, $\mathcal{U}_{\theta}$ is a learned function that progressively refines $\bm{x}_{t}$ towards the target distribution by reversing the noising process defined by Equation (1).
While iteratively applying $\mathcal{U}_{\theta}$ enables sampling from the posterior distribution, it does not natively incorporate externally defined constraints, or even implicit ones, given the stochastic nature of the process. By introducing a constraint set $\mathbf{C}$ and enforcing $\bm{x}_{t}\in\mathbf{C}$ at each step, CDD extends the reverse process to sample from these constrained subdistributions. This is operationalized by the introduction of a projection operator:
$\bm{x}_{s}=\mathcal{P}_{\mathbf{C}}\bigl{(}\mathcal{U}_{\theta}(\bm{x}_{t})\bigr{)}$
computing the “nearest sample” that falls within the feasible subdistribution according to a distance metric defined over probability distributions. CDD formulates this projection as a constrained optimization problem that minimizes the Kullback–Leibler (KL) divergence between the projected and original probability distributions, ensuring minimal deviation, while enforcing constraint adherence. Given the model’s predicted probabilities ${\bm{x}_{t}}\in\mathbb{R}^{L\times V}$, the projection is defined as:
$\mathcal{P}_{\mathbf{C}}(\bm{x}_{t})=\arg\min_{\bm{y}}D_{\mathrm{KL}}\left(\mathcal{U}_{\theta}(\bm{x}_{t})\|\bm{y}\right)$ (2a)
[…]
where $\mathbf{C}$ represents the constraint set governing valid outputs. The projection step is formulated as a subproblem of the sampling procedure, which can be solved using gradient-based methods, as we discuss in section 4.2.
Importantly, note that while the process relies on metrics computed at the token probability distributions, valid outputs are assessed at the sequence level. Indeed, the constraints are formulated such that the $\arg\max$ of the projected sequence $\bm{y}$ (referred to as $\bm{y}^{\star}$), must satisfy sentence level criteria $\bm{y}^{\star}\in\mathbf{C}$ [^1] We assume a greedy decoding scheme, hence the use of $\arg\max$, as is standard to current diffusion-based language models. However, this use of a non-differentiable $\arg\max$ operation poses challenges to the resolution of the projection, which relies on gradient-based optimization frameworks.
Differentiating over $\mathbf{\arg\max}$ operators.
To address the non-differentiability of the $\arg\max$ function, we apply a Gumbel-Softmax relaxation $\tilde{\phi}$, which approximates discrete token probabilities with continuous values [38] as
$\tilde{\phi}(\bm{x}_{t}^{i})(v)=\frac{\exp\!\Bigl{(}\frac{\log\bm{x}_{t}^{i}(v)+g_{v}}{\text{T}_{\text{sample}}}\Bigr{)}}{\sum_{v^{\prime}=1}^{V}\exp\!\Bigl{(}\frac{\log\bm{x}_{t}^{i}(v^{\prime})+g_{v^{\prime}}}{\text{T}_{\text{sample}}}\Bigr{)}},$
where $\bm{x}_{t}^{i}(v)$ is the probability of token $v$ in the vocabulary, $g_{v}$ is drawn from the Gumbel$(0,1)^{1}$ distribution for token $v$, and $\text{T}_{\text{sample}}>0$ is the temperature parameter controlling the smoothness of the output. This allows gradients to be propagated to solve the projection step while closely approximating the discretized $\arg\max$ operation.
4.2 Enforcing Constraints
Consider a generic constraint on a sequence $\bm{x}$ defined via a measurable property given by a function $g(\bm{x})$. For instance, $g(\cdot)$ might evaluate sentence toxicity (see Section 5.1), count the number of words occurring in a string (see Section 5.2), or even serve as a black-box function computed by an external routine, as in our molecular sequence generation application (see Section 5.3). To guide the projection operations, we require a measure of constraint violation that can parametrize the update rule of the gradient-based projection method. For ease of presentation, we assume $g$ is non-negative and we express the constraint in the form $g(\bm{x})<\tau$, where $\tau\geq 0$ represents an acceptable threshold that must not be exceeded.
To quantify by how much a given sequence violates the constraint, we define,
$\Delta g(\tilde{\phi}(\bm{x}_{t}))=\max\left(0,g(\tilde{\phi}(\bm{x}_{t}))-\tau\right),$
where $\tilde{\phi}(\bm{x}_{t})$ is a projection or transformation that ensures $g$ is well-defined on $\bm{x}$. If $g(\tilde{\phi}(\bm{x}_{t}))\leq\tau$, the differentiable relaxation of the constraint is satisfied and $\Delta g$ is zero; otherwise, $\Delta g$ measures the violation.
Since $\Delta g$ is typically non-linear and difficult to impose directly, we adopt a method inspired by the augmented Lagrangian approach [39], where the constraints are incorporated into the objective of a minimizer via Lagrange multipliers $\lambda$ and a quadratic penalty term $\mu$. Let $\mathcal{U}_{\theta}(\bm{x}_{t})$ be the probability distribution after the denoising step at diffusion time $t$. We introduce a projected distribution $\bm{y}$, which is iteratively updated to reduce the constraint violations (measured by the score $\Delta g$) while remaining close (in KL-divergence) to $\mathcal{U}_{\theta}(\bm{x}_{t})$. Concretely the agumented Lagrangian dual function is defined as:
$\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu)=\;$ $D_{\mathrm{KL}}\left(\mathcal{U}_{\theta}(\bm{x}_{t})\|\bm{y}\right)$
+\;+ $\lambda\Delta g(\tilde{\phi}(\bm{y}))$
+\;+ $\frac{\mu}{2}\Delta g(\tilde{\phi}(\bm{y}))^{2}$
where $\lambda$ is a Lagrange multiplier, and $\mu$ is a quadratic penalty term. When using a Lagrangian function, the optimization problem becomes
$\arg\min_{\bm{y}}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu),$
and is a lower bound of the original projection operator (2). To obtain the strongest Lagrangian relaxation of the projection, the Lagrangian dual can be used to find the best Lagrangian multiplies $\lambda$ and penalty terms $\mu$, i.e.,
$\arg\max_{\lambda,\mu}\Bigl{(}\arg\min_{\bm{y}}\bigl{(}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu)\bigr{)}\Bigr{)}.$ (3)
For various classes of problems, the Lagrangian dual is a strong approximation of the original problem (our projection into the constraint space).
Algorithm 1 Augmented Lagrangian Method
1:Init: $\lambda,\mu,\eta,\alpha,\delta$
2:Input: probability distribution $\mathcal{U}_{\theta}(\bm{x}_{t})$
3:$\bm{y}=\mathcal{U}_{\theta}(\bm{x}_{t})$
4:while $\Delta g(\bm{y}^{*})<\delta$ do
5:     for $j=1$ to max_inner_iter do
6:         $\mathcal{L}_{KL}\leftarrow D_{\mathrm{KL}}(\mathcal{U}_{\theta}(\bm{x}_{t})\|\bm{y})$
7:         $\mathcal{L}_{\text{ALM}}\leftarrow\mathcal{L}_{KL}+\lambda\,\Delta g(\tilde{\phi}(\bm{y}))+\tfrac{\mu}{2}\,\Delta g(\tilde{\phi}(\bm{y}))^{2}$
8:         ${\bm{y}}\leftarrow{\bm{y}}-\eta\,\nabla_{\bm{y}}\mathcal{L}_{\text{ALM}}$
9:     end for
10:     $\lambda\leftarrow\lambda+\mu\,\Delta g(\bm{y}^{*})$
11:     $\mu\leftarrow\min\bigl{(}\alpha\mu,\,\mu_{\text{max}}\bigr{)}$
12:end while
13:$\bm{x}_{s}=\bm{y}$
14:return $\bm{x}_{s}$
The optimization of (3) proceeds iteratively, following a gradient-based update on $\bm{y}$ while dynamically adjusting the Lagrange multiplier $\lambda$ and penalty coefficient $\mu$. Specifically, we perform the following updates [40]:
$\bm{y}$ $\leftarrow\bm{y}-\eta\nabla_{\bm{y}}\mathcal{L}_{\mathrm{ALM}}(\bm{y},\mathcal{U}_{\theta}(\bm{x}_{t});\lambda,\mu)$ (4a)
$\lambda$ $\leftarrow\lambda+\mu\Delta g(\bm{y}^{\star}),$ (4b)
$\mu$ $\leftarrow\min(\alpha\mu,\mu_{\max}),$ (4c)
where $\bm{y}^{\star}=\arg\max(\bm{y})$ is the discrete sequence implied by the distribution $\bm{y}$, $\eta$ is the gradient step size, $\alpha>1$ is a scaling factor that progressively increases $\mu$ over iterations, and $\mu_{\max}$ is an upper bound on the penalty term. These updates drive $\bm{y}$ as close as possible to satisfy $\Delta g(\tilde{\phi}(\bm{y}^{\star}))\leq\tau$ while also ensuring it remains close to the original denoised distribution ${\cal U}_{\theta}(\bm{x}_{t})$. We provide pseudocode for our implementation in Algorithm 1.
As we will show in the next section, practically these constraints can also be implemented as surrogate models (e.g., a classifier) that can be used to provide a continuous score, allowing for smooth gradient-base updates, as shown above. Additionally, we will also show that for token-level or logical constraints, such as enforcing the presence or absence of specific (combination of) tokens, $g$ can be an identity function, leading to sharper gradients and faster convergence in the iterative steps of Equation (4).
5 Experiments
To empirically demonstrate the advantage provided by CDD, we compare our method against similarly sized autoregressive and diffusion-based language models. Specifically, we benchmark our performance on constrained text generation against autoregressive baselines: GPT-2 and Llama 3.2 [41, 42], and discrete diffusion baselines MDLM and UDLM [16, 17]. We intentionally select models of similar size to, and up to one magnitude larger than, our discrete diffusion model, noting that we expect the performance of both classes of models to scale similarly with size. As shown by Schiff et al. [17], MDLM outperforms UDLM on the tasks described in Sections 5.1 and 5.2, and therefore we adopt this stronger baseline for comparison in these sections and utilize both in Section 5.3. For this reason, we use MDLM as the base discrete diffusion model for Sections 5.1 and 5.2, and UDLM for Section 5.3.
We evaluate on three constrained language generation tasks:
  1. 1.
    Toxicity Mitigation: Given a prompt from RealToxicityPrompts [43], the model is tasked with completing the passage with the goal of providing non-toxic responses.
  2. 2.
    Instruction Following: We present lexical constraints in an instruction-tuning format and ask the model to generate outputs that adhere to these specifications.
  3. 3.
    Molecule Generation: Our language model, trained on the molecular grammar ‘SMILES’ [44], is tasked to generate novel molecule sequences or sequences which satisfy certain properties.
5.1 Toxicity Mitigation
While LLMs have demonstrated remarkable capabilities in generating human like text, a byproduct of data driven generation is the inadvertent creation of harmful, offense and dangerous text. Despite undergoing alignment finetuning, current state-of-the-art models still frequently generate harmful output, and while post-proccesing methods can be effective in moderating generations, they have still been shown to generate toxic output as these can be circumvented by adversarial or malicious prompts.
For this experiment, we generate output conditioned on prompts from the RealToxicityPrompts dataset [43]. We select prefixes with toxicity score above 0.5, in order to test how these models perform when given toxic inputs, and perplexity below 30.0, to avoid incoherent generations. To model the constraints, we finetune a GPT-Neo [45] model on the Jigsaw Toxic Comment Classification Challenge [^2] https://www.kaggle.com/c/jigsaw-toxic-comment- classification-challenge.
Figure 3 presents a comprehensive evaluation on model performance for toxicity mitigation, where the objective is to satisfy some toxicity threshold based on a surrogate model while maintaining competitive perplexity and language fluency. In this experiment, we assess several models architectures, baselines, and different toxicity thresholds. We report perplexity (PPL) at various percentiles (mean, 25th, 50th, and 75th), and violation rates at toxicity thresholds $\tau=0.25$, $\tau=0.50$, and $\tau=0.75$. All perplexity calculations are computed with respect to GPT2 (774M).
Among the baselines, GPT2 (124M) and Llama (1B) achieve the lowest perplexity score, however, they frequently produce toxicity content resulting in high violations. Although GPT2 + PPLM (345M) [30] is able to decrease the violations for all toxicity thresholds, it is unable to consistently prevent toxic generations and suffers from an increase in perplexity. MDLM (110M) exhibits higher perplexity than GPT2, with a median PPL of 39.75, and although toxicity violation rates are moderately improved relative to GPT2, they remain non-negligible. In contrast, our Constrained Discrete Diffusion (CDD) provides perfect constraint satisfaction at each toxicity thresholds, maintaining reasonable perplexity scores.
Toxicity Model Size PPL Viol (%)
Mean 25th 50th 75th $\tau=0.25$ $\tau=0.50$ $\tau=0.75$
GPT2 124M 19.143 14.57 17.75 21.61 36.3 23.8 16.2
[…]
Llama 3.2 1B 15.66 11.52 14.58 17.95 34.9 27.8 23.1
MDLM 110M 46.72 26.57 39.75 57.55 32.1 23.2 17.2
$\textbf{CDD}_{\tau=0.25}$ (Ours) 110M 73.68 30.48 47.60 74.07 0.0 0.0 0.0
$\textbf{CDD}_{\tau=0.50}$ (Ours) 110M 60.49 29.14 45.64 68.10 0.0 0.0
$\textbf{CDD}_{\tau=0.75}$ (Ours) 110M 56.66 28.35 44.31 65.19 0.0
Figure 3: Results for various toxicity thresholds and PPL percentiles.
These results illustrate the trade-off between perplexity and toxicity mitigation. Baseline models such as GPT2 and Llama 3.2 achieve high fluency but at the cost of increased toxic outputs, whereas methods like GPT2 with PPLM provides neither strict avoidance of toxic generation nor the highest degrees of fluency. Our method, by contrast, provides a mechanism to establish a strict threshold on toxicity with respect to an external evaluator.
These findings are important as they show how, using projection-based control strategies, it is possible to achieve low toxicity levels at desired thresholds.
5.2 Instruction Following
Current state-of-the-art language models often struggle to satisfy seemingly simple generation requests from the user, and while instruction tuning has been shown effective in reducing the frequency that LLMs deviate from desired behaviors, these models still struggle to consistently satisfy user-defined constraints.
We evaluate CDD against such models using two types of constraints. First, we assess the ability of CDD to accurately answer how many letters are in a word, such as ‘How many R’s are in Strawberry?’. We additionally test lexical constraints based of the COLLIE dataset [46], prompting the models to generate word-indexed constrained sentences,
Model Size PPL Viol (%)
Counting GPT-2 (zero-shot) 124M 19.2 ± 0.8 100
GPT-2 (instruction-tuned) 124M 23.4 ± 1.3 5.0
Llama 3.2 (zero-shot) 1B 34.3 ± 3.8 100
Llama 3.2 (instruction-tuned) 1B 30.7 ± 2.9 9.1
MDLM 169M 16.9 ± 1.0 54.5
CDD (Ours) 169M 17.9 ± 1.1 0.0
Lexical GPT-2 (zero-shot) 124M 26.6 ± 4.2 97.6
GPT-2 (instruction-tuned) 124M 62.2 ± 2.9 96.4
Llama 3.2 (zero-shot) 1B 11.3 ± 2.0 97.2
Llama 3.2 (instruction-tuned) 1B 66.5 ± 6.1 93.6
MDLM 169M 21.9 ± 2.0 97.5
CDD (Ours) 169M 26.6 ± 4.1 0.0
Table 1: Results for functional constraints. PPL is assessed using GPT2 (774M). Bold and underlined values mark the best and second-best, respectively.
such as ‘Generate a sentence where the third-to-last word is mankind’. For each constraint we formulate a projection operator to impose hard constraints on the generation, pushing the samples back onto the constraint set.
5.2.1 Counting Constraints
As illustrated in the top of Table 1, unconstrained GPT-2 and Llama 3.2 fail to satisfy any letter-count requirement. Instruction tuning helps, reducing violation rates to 5.0% (GPT-2) and 9.1% (Llama 3.2), but still leaves a notable error margin. Although MDLM achieves the lowest perplexity, over half of its outputs violate the constraint. By contrast, CDD fully eliminates violations while preserving near-best fluency (17.9 perplexity), demonstrating that its constrained sampling process provides strict compliance without sacrificing language quality.
5.2.2 Lexical Constraints
Results are reported at the bottom of Table 1. Zero-shot GPT-2 and Llama 3.2 show similarly high violation rates (about 97%), and instruction tuning only minimally improves these figures. MDLM again delivers relatively low perplexity (21.9) but fails to address the precise positional requirements (97.5% violations). In stark contrast, CDD ensures perfect adherence, with only a moderate increase in perplexity, thus confirming the effectiveness of its projection mechanism for enforcing hard lexical constraints.
Taken together, these experiments reveal a fundamental limitation of conventional and instruction-tuned language models in scenarios requiring strict constraints. Conversely, CDD effectively produces outputs that satisfy both counting and lexical requirements by leveraging a projection operator to impose these constraints during generation.
Molecules (Synthetic Accessibility) Model Valid Novel QED Viol (%)
[]delimited-[][$\uparrow$][ ↑ ] []delimited-[][$\uparrow$][ ↑ ] []delimited-[][$\uparrow$][ ↑ ] $\tau=3.0$ $\tau=3.5$ $\tau=4.0$ $\tau=4.5$
AR 1023 11 0.46 91.6 78.4 62.3 42.0
$\text{AR}_{\text{FUDGE}\>\gamma=7}$ 925 13 0.48 11.1 9.8 9.6 9.2
MDLM 596 271 0.45 85.9 73.7 61.1 44.0
$\text{MDLM}_{\text{D-CFG}\>\gamma=3}$ 772 53 0.41 87.8 73.9 54.2 22.5
$\text{MDLM}_{\text{D-CBG}\>\gamma=3}$ 436 21 0.37 50.5 48.6 46.1 44.7
UDLM 895 345 0.47 89.4 88.0 58.1 37.8
$\text{UDLM}_{\text{D-CFG}\>\gamma=5}$ 850 69 0.47 80.6 58.6 35.9 13.9
$\text{UDLM}_{\text{D-CBG}\>\gamma=10}$ 896 374 0.47 90.1 77.8 58.6 37.7
$\textbf{CDD}_{\tau=3.0}$ (Ours) 353 36 0.63 0.0 0.0 0.0 0.0
$\textbf{CDD}_{\tau=3.5}$ (Ours) 863 91 0.62 - 0.0 0.0 0.0
$\textbf{CDD}_{\tau=4.0}$ (Ours) 936 108 0.61 - - 0.0 0.0
$\textbf{CDD}_{\tau=4.5}$ (Ours) 938 121 0.58 - - - 0.0
Figure 4: Results for synthetic accessibility constrained molecule generation constraints (left) and illustration of projection operation (right). QED and constraint violations are reported for only valid molecules. Bold and underlined values mark the best and second-best, respectively.
5.3 Molecular Generation
In this experiment we use UDLM to generate SMILES strings [44], a linear representation of a molecule’s structure with the vocabulary consisting of a finite set of symbols, and a grammar dictating molecular validity. Recent advances in generative modeling have enabled the design of molecular sequences by leveraging techniques from natural language processing. Despite their impressive ability to optimize for specific chemical properties, these models often generate molecules that fall short of practical requirements, either by producing compounds that are difficult to synthesize or by closely mimicking existing structures. Furthermore, generated molecules often fail to maximize qualitative properties such as drug-likeness (QED) [47] that are critical to the practicality of the generations and used as a qualitative metric for the experiments discussed in this section. This gap between in silico performance and laboratory feasibility has become a major hurdle in advancing molecular discovery.
To address this challenge, our framework incorporates two distinct constraint mechanisms. The first focuses on synthetic feasibility: by employing a surrogate scoring function, we ensure that every generated molecule meets a strict threshold for ease of synthesis, effectively prevent the generation of candidates that would be impractical to produce [48]. The second constraint targets molecular originality by rigorously projecting sequences that fall within a set of
Molecules (Novelty) Model Size Valid & Novel QED Viol (%)
No Guidance
AR 92M 11 0.41 98.93
MDLM 92M 271 0.45 54.53
UDLM 92M 345 0.46 61.45
CDD 92M 511 0.45 0.00
CFG
$\text{AR}_{\text{D-CFG}\>\gamma=3}^{\dagger}$ 92M 79 0.60 91.61
$\text{MDLM}_{\text{D-CFG}\>\gamma=3}^{\dagger}$ 92M 96 0.60 69.82
$\text{UDLM}_{\text{D-CFG}\>\gamma=5}^{\dagger}$ 92M 64 0.62 93.69
$\textbf{CDD}_{\text{D-CFG}\>\gamma=5}(\text{Ours})$ 92M 251 0.60 0.00
CBG
$\text{AR}_{\text{FUDGE}\>\gamma=7}^{\dagger}$ 92M 53 0.61 94.28
$\text{MDLM}_{\text{D-CBG}\>\gamma=3}^{\dagger}$ 92M 117 0.58 72.08
$\text{UDLM}_{\text{D-CBG}\>\gamma=10}^{\dagger}$ 92M 64 0.61 93.59
$\textbf{CDD}_{\text{D-CBG}\>\gamma=10}(\text{Ours})$ 92M 355 0.59 0.00
Table 2: Results for novelty projection with and without QED guidance. Violation represents percentage of valid, but not novel molecule generations. QED is reported for only novel molecules. Results denoted with are as reported by Schiff et al. [17]. Bold and underlined values mark the best and second-best, respectively.
known compounds. For each constraint, we implement a dedicated projection operator that imposes hard limits on the generation process, resulting in final outputs that adhere precisely to our desired standards of synthetic accessibility or novelty. The ability to impose such constraints marks a crucial step toward the development of autonomous, LLM-driven laboratories for molecular innovation.
5.3.1 Synthetic Accessibility Constraint
While our molecular generation framework employs a surrogate model to optimize for synthetic accessibility during training, the actual assessment of accessibility violations is conducted by a separate, black-box external model. This external evaluation rigorously measures the degree to which generated molecules comply with the synthetic accessibility criteria, using a series of thresholds ($\tau=3.0$, $3.5$, $4.0$, and $4.5$). The results, summarized in Figure 4, confirm that CDD achieves perfect compliance across all thresholds. Across all thresholds, the CDD model yields a 0% violation rate, despite generating a competitive number of valid molecules and exhibiting the highest drug-likeness scores (referred to in the table as QED). These findings underscore the effectiveness of our hard constraint projection operators in ensuring that every candidate molecule not only scores well in property optimization but also adheres strictly to synthetic feasibility requirements as determined by an independent, external standard.
5.3.2 Novelty Constraint
Novelty in molecule generation refers to the model’s ability to produce new chemical structures that were not explicitly contained in its training set. In the context of generative modeling for drug design or other chemistry applications, novelty is often an important objective for the reasons of chemical space exploration and practical relevance. In real-world scenarios such as designing new drugs, researchers already have access to known molecules, thus, generating truly novel structures is crucial for discovering new compounds or structures.
In this experiment, we generate SMILES strings constraining all valid generations to be novel. We use baslines with no QED guidance alongside CFG and CBG to evaluate this methods impact on a desirable guidance improved metric. During the denoising step, if the candidate sequence is in an external database, we run a projection to generate the smallest possible change in probability space resulting in a novel molecule, effectively approximating the gradient step to this non-differentiable novelty indicator function. By applying this approach at every diffusion step, we ensure the model only produces new, previously unseen molecules, while simultaneously taking advantage of the creativity of the underlying discrete diffusion model.
As shown in Table 2, the CBG setting yields a 203.4% increase in novel molecule generation, while the CFG setting shows a 161.4% increase. Even without guidance, the method still produces 48.1% more novel molecules, with only a minimal reduction in the QED score.
These results demonstrate CDD’s applicability for scientific applications. Ensuring novelty is imperative for molecule discovery, while chemical property adherence is crucial for real world applications in drug-design or materials discovery.
6 Limitations
There are two key limitations that are currently faced in the development of constrained discrete models. The first regards the size of currently available discrete diffusion models. While current discrete diffusion models provide comparable performance in fluency and perplexity to similarly sized autoregressive models, including state-of-the-art models such as Llama 3.2 1B, their generative capabilities fall short of large-scale language models. Larger scale, closed-source discrete diffusion models have been developed and boasted impressive performance [37], but larger open-source models are yet to be provided to the research community. It is expected that the release of such larger models will further boost the performance of our proposed methods.
The second key limitation concerns an inherent byproduct of constraint imposition: increased computational overhead. While unconstrained discrete diffusion language models have been shown to provide significant speed-ups over autoregressive language models due to the parallelization of token generation [18, 16, 49], the inclusion of constraints within the generation process may limit this benefit. Nevertheless, this trade-off is often justified, as imposing constraints is essential in scenarios where correctness outweighs generation speed, such as ensuring safety-critical outputs or supporting rigorous scientific exploration. In such contexts, achieving accurate and reliable outcomes is prioritized over computational efficiency.
7 Conclusion
In this paper we presented Constrained Discrete Diffusion (CDD), a novel method for imposing both structured and unstructured constraints during language generation. CDD relies on the adoption of discrete diffusion models which enable the consideration of entire sequences when imposing constraints during the generation process. Offering a key departure from current paradigms for controllable text generation, CDD integrates differentiable optimization methods into the diffusion reverse process. This integration, governed by a rigorous mathematical optimization framework, provides precise control for natural language generation. The method is empirically validated on (i) toxicity mitigation and harmful content prevention, (ii) imposition of lexical constraints at character and sequence levels, and (iii) generation of novel molecular sequences that adhere to specific properties for drug discovery, reporting state-of-the-art results for controllable generation in both terms of constraint adherence and quality metrics.
Acknowledgments
This work was partially supported by NSF grants CAREER-2401285 and RI-2334936 and UVA’s National Security Data & Policy Institute, ODNI Contracting Activity #2024-24070100001 Its view and conclusions reflect those of the authors only. The work has also been supported by LLNL under Contract DE-AC52- 07NA27344 and by the LLNL-LDRD Program under Project No. 24-ERD-010 and 24-SI-008. This manuscript has been co-authored by Lawrence Livermore National Security, LLC under Contract No. DE-AC52-07NA27344 with the U.S. Department of Energy. The United States Government retains, and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes.
Appendix A Expanded Related Work
Constrained Decoding Algorithms.
A prominent strategy in controllable text generation is to employ constrained decoding algorithms, which incrementally parse outputs under specified formatting or structure requirements. Such approaches have been extensively used in tasks like code generation [50], semantic parsing [51], SQL queries [52], and grammar-constrained text generation [20]. These methods typically prune the vocabulary at each decoding step to only those tokens that comply with the target representation’s syntax or semantics. For instance, Geng et al. confine the logit distribution to tokens valid within a formal grammar, then employ a completion engine to extend partial sequences while maintaining grammatical consistency. Meanwhile, approaches that treat the original language model as a black box resort to external constraints. For example, [21] harnesses a secondary model to parse and restrict outputs, an option that becomes necessary when direct access to the logits is unavailable. However, these workarounds may suffer performance penalties due to their indirect control over the primary model’s probability distribution.
Beyond syntax-focused constraints, lexical constraints have also been extensively studied. Lu et al. [22] devise Neurologic decoding, which encourages the presence or absence of specific words in generated text via a modified beam search. They later improve adherence to these constraints by employing an A* search variant [23]. Similar frameworks combine toxicity classifiers with beam search to steer generation toward more benign text [30, 53]. Although these methods can effectively address a range of constraints, they generally rely on augmented sampling heuristics or additional classifiers rather than natively integrating constraints into the generation process.
Natural Language Verification with Formal Logic.
Another line of work has examined mapping natural language into logical representations to ensure factual and logical consistency [24, 25, 26]. By encoding linguistic inputs in a formal system, these methods can detect inconsistencies or factual errors and thereby provide a measure of control over generated outputs. For instance, Olausson et al., Jiang et al., and Zhang et al. convert natural language statements into first-order logic (FOL) and leverage theorem proving to verify factual correctness. However, FOL struggles to capture complex temporal or sequential dependencies, limiting its applicability to planning or process-oriented tasks. Similarly, other work uses entailment models to verify that a conclusion logically follows from a set of premises [57], though these methods generally stop short of actively generating text that satisfies domain-specific constraints.
Agent Constraint Satisfaction.
In robotics and embodied AI, natural language planning frequently incorporates specialized external modules to handle feasibility constraints. For example, Chen et al. [27] and Lin et al. [28] defer to domain-specific motion planners to generate viable trajectories, embedding constraint satisfaction directly into the underlying optimization routines. Although these solutions can integrate strong, domain-aware constraints, they do so by delegating control to highly specialized components. On the discrete side, Yang et al. [58] achieves constraint satisfaction by running a validation module at each step of plan construction and discarding infeasible actions through rejection sampling. While effective in certain settings, such methods typically require domain-specific expertise and do not uniformly integrate constraint enforcement into the generative model itself.
Appendix B Experimental Setup
B.1 Toxicity Mitigation
Surrogate Model.
As the surrogate model for toxicity task, we use a GPT-Neo (1.3B) model, adapted for binary classification. We finetune this model on the Jigsaw toxicity dataset which includes multiple toxicity-related labels such as toxic, severe toxic, insult, etc. We consolidates these columns into a single binary target (toxic vs. non-toxic).
B.2 Counting Constraints
Instruction Tuning Dataset.
We manually construct a dataset of letter counting examples in an instruction-tuning format. We use random words and select a random character from these words 80% of the time. The other 20% of the time we select a letter at random, which may or may not be contained in the word, to represent cases where the correct answer is zero.
Projection Operator.
Our implementation of this projection operator involves integrating an external module to count the number of times a specific letter appears in a given word. First, we leverage the external module to obtain the accurate count, ensuring that the correct numerical value is determined. Then, during output generation, we adjust the probability distribution over the numerical tokens so that the token corresponding to this count is assigned the highest probability, while the probabilities of other numerical tokens are appropriately diminished. This approach guarantees that when the model is queried about the frequency of a letter in a word, it outputs the correct numerical token with the highest weight in the distribution, maintaining consistency and precision in the final answer.
B.3 Lexical Constraints
Instruction Tuning Dataset.
We finetune models on a dataset of examples constructed from evaluation of GPT-4 on the COLLIE dataset provided by Yao et al. [46]. As GPT-4 performs well on these examples, we apply knowledge distillation by compiling these generations into an instruction tuning dataset. The dataset can be directly constructed from the logs included in the code base from Yao et al.
Projection Operator.
Our implementation of the projection operator begins by parsing the decoded string to locate the position corresponding to the $n_{th}$ word. Because token indices do not necessarily match word indices, owing to the nuances of subword tokenization, we perform an alignment procedure that accurately maps word positions to their respective token indices. Once the index in the sequence corresponding to the $n_{th}$ word is identified, we apply a closed-form gradient update that adjusts the probability distribution over the vocabulary. This update maximizes the probability of the correct token at the specified position while ensuring that the overall distribution remains valid (i.e., all probabilities sum to one).
B.4 Molecular Synthesis
Training Dataset.
For our molecule generation experiments, we utilize the QM9 dataset [59], as used by Schiff et al. [17]. This dataset comprises approximately 133,885 small organic molecules, each encoded as a SMILES string [44], which compactly represents the molecular structure through a sequence of characters. The SMILES representation facilitates discrete modeling by enabling the treatment of molecular generation as a sequence prediction task, making it particularly amenable to discrete diffusion approaches. By leveraging QM9, we can rigorously evaluate the performance of our generative models on tasks that require both the preservation of chemical validity and the precise control of molecular properties, aligning with established protocols in the literature.
Figure 5: Synthetic Accessibility score distributions for CDD versus competing baselines. Importantly, CDD is the only model that never violates the specified toxicity thresholds.
Synthetic Accessibility Surrogate Model.
We train our surrogate model on the QM9 dataset [59] but manually label the training data with with RDKIT’s synthetic accessibility score [60]. We finetune GPT2 (124M) to act as this surrogate model and directly output a score $s\in[0,10]$.
Novelty Projection Operator.
In order to project the molecule sequences into a novel set, we apply a best-first search (BFS) at the probability distribution level. We begin by defining a flip cost for forcing a token to become the new argmax; this cost is the difference between the current top token’s probability and the candidate token’s probability, or zero if the candidate is already top-ranked. To find a novel sequence at minimal total flip cost, we start with an empty sequence and expand it position by position, accumulating flip costs in a priority queue. Once a full sequence is constructed, we decode it and check if it is absent from our existing dataset. If it is indeed novel, we finalize the sequence, shift the probability distribution so that each selected token becomes the definitive argmax, and insert the resulting sequence into the dataset to prevent re-generation. This procedure systematically maintains high-likelihood sequences while avoiding those already present, terminating for each sample as soon as it finds a suitable novel result before proceeding to the next sample.